1. Fixed Cost is 100, calculate TC, AC, MC, AFC and AVC 

Qty. TVC

1. 500

2 600

3 850

4 980

5 1020


1
Expert's answer
2021-05-27T19:07:11-0400

a) Total cost can be found as follows:


"TC_1=TFC_1+TVC_1=\$100+\$500=\$600,""TC_2=TFC_2+TVC_2=\$100+\$600=\$700,""TC_3=TFC_3+TVC_3=\$100+\$850=\$950,""TC_4=TFC_4+TVC_4=\$100+\$980=\$1080,""TC_5=TFC_5+TVC_5=\$100+\$1020=\$1120."

b) Average cost can be found as follows:


"AC_1=\dfrac{TC_1}{Q_1}=\dfrac{\$600}{1}=\$600,""AC_2=\dfrac{TC_2}{Q_2}=\dfrac{\$700}{2}=\$350,""AC_3=\dfrac{TC_3}{Q_3}=\dfrac{\$950}{3}=\$316.6,""AC_4=\dfrac{TC_4}{Q_4}=\dfrac{\$1080}{4}=\$270,""AC_5=\dfrac{TC_5}{Q_5}=\dfrac{\$1120}{5}=\$224."

c) Marginal cost can be calculated as follows:


"MC_1=\dfrac{TC_2-TC_1}{Q_2-Q_1}=\dfrac{\$700-\$600}{2-1}=\$100,""MC_2=\dfrac{TC_3-TC_2}{Q_3-Q_2}=\dfrac{\$950-\$700}{3-2}=\$250,""MC_3=\dfrac{TC_4-TC_3}{Q_4-Q_3}=\dfrac{\$1080-\$950}{4-3}=\$130,""MC_4=\dfrac{TC_5-TC_4}{Q_5-Q_4}=\dfrac{\$1120-\$1080}{5-4}=\$40."

d) Average fixed cost can be found as follows:


"AFC_1=\dfrac{TFC}{Q_1}=\dfrac{\$100}{1}=\$100,""AFC_2=\dfrac{TFC}{Q_2}=\dfrac{\$100}{2}=\$50,""AFC_3=\dfrac{TFC}{Q_3}=\dfrac{\$100}{3}=\$33.3,""AFC_4=\dfrac{TFC}{Q_4}=\dfrac{\$100}{4}=\$25,""AFC_5=\dfrac{TFC}{Q_5}=\dfrac{\$100}{5}=\$20."

e) Average variable cost can be found as follows:

"AVC_1=\dfrac{TVC_1}{Q_1}=\dfrac{\$500}{1}=\$500,""AVC_2=\dfrac{TVC_2}{Q_2}=\dfrac{\$600}{2}=\$300,""AVC_3=\dfrac{TVC_3}{Q_3}=\dfrac{\$850}{3}=\$283.3,""AVC_4=\dfrac{TVC_4}{Q_4}=\dfrac{\$980}{4}=\$245,""AVC_5=\dfrac{TVC_5}{Q_5}=\dfrac{\$1020}{5}=\$204."

Finally, we get the following table: